OpenStudy (noahbred):

What is the correct reason for statement 2? Prove: 4 + 3x + (x ∙ 2) = 5x + 4 Statements Reasons 1. 4 + 3x + (x ∙ 2) 1. Given 2. 4 + 3x + 2x 2. 3. 4 + 5x 3. Addition 4. 5x + 4 4. Commutative Property A. Associative Property B. Commutative Property C. Multiplication D. Addition @lightn1ng

OpenStudy (lightn1ng):

kk

OpenStudy (danjs):

which property says the order of multiplication doesnt matter x * y = y * x

OpenStudy (anonymous):

@noahbred, can you tell me what each of the answers A-D mean?

OpenStudy (danjs):

going from step 1 to 2, they just changed x*2 to 2*x

OpenStudy (lightn1ng):

multiplication because he multiplied whats in the brackets to simplify the equation. So the answer is C.

OpenStudy (lightn1ng):

This is true ^ however they did not switch them. He just multiplied whats in the brackets.

OpenStudy (danjs):

they are always multiplied, x times 2, and then 2 times x

OpenStudy (danjs):

2x is 2 times x

OpenStudy (danjs):

change in the order of multiplication... like 5*3 = 3*5 for instance

OpenStudy (anonymous):

@DanJS, I think you made a mistake in what you said before (or I just misunderstood!). Both addition and multiplication have associative and commutative properties: Associative Multiplication: a*(b*c)=(a*b)*c Commutative Multiplication: a*b=b*a Associative Addition: a+(b+c)=(a+b)+c Commutative Addition: a+b=b+a

OpenStudy (danjs):

commuinative , sorry just fixed it

OpenStudy (danjs):

communative -- -order of it associative -- which pairs first

OpenStudy (danjs):

sorry

OpenStudy (danjs):

the other one shown had a 4*(2x) ---> (4*2)*x both distributive and associative are choices, would distributive just remove the parenthesis, and that is really associative